The probability that a randomly selected teenager studied at least once during a week was only 0.67. What is the probability that less than half of the students in your study group of 10 have studied in the last week? (Round your answer to two decimal places.)
Answer #1

X ~ Bin ( n , p)

Where n = 10 , p = 0.67

P(X) = nCx * px * ( 1 - p)n-x

P(X < 5) = ?

P(X < 5) = P(X <= 4)

= P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3) + P(X = 4)

= 10C0 * 0.670 * ( 1 - 0.67)10 + 10C1 * 0.671 * ( 1 - 0.67)9 + 10C2 * 0.672 * ( 1 - 0.67)8 +

10C3 * 0.673 * ( 1 - 0.67)7 + 10C4 * 0.674 * ( 1 - 0.67)6

= 0.0732

= 0.07 (Rounded to two decimal places)